Ram said to Mohan,”I am twice ass old as you were when I was old as you.” The sum of their present ages is 49 years .Then find their ages in years are.

Grade:11

1 Answers

ganesh ahire
5 years ago
let ram’s age be R
and mohan’s  be M
as given
R+M=49                                                                                             (1)
now as per condition 
ram is twice mohan’s age when he himself was of Mohans present age  
means                                      
this condition tells ram is of  M years at that time
and
we need to calculate  how much the both have grown up ...
Which is difference  of rams present and conditional age
let it be D,
&    D= R-M
now as per condition ram is twice mohans present age subtracted by D
therefore   
R/2 = M-(R-M)
R/2 = 2M-R
3R/4= M                                                                 (2)
put the result (2)  in  1 to get linear equation in one variable 
3R/4+R=49
7R/4=49
R= 28
and so by (1)           M= 21
so ram is 28 and mohan 21